Home Canning Products common stock sells for $41.00 a share and has a market rate of return of 15.9 percent. The company just paid an annual dividend of $2.95 per share. What is the dividend growth rate?
Answer #1

This question requires application of constant growth dividend discount model according to which: Divi Po T-9 Po = Price of S

Div1 = $2.95 * (1 + g)

41 = \frac{2.95 * (1 + g)}{0.159 - g}

6.519 - 41g = 2.95 + 2.95g

3.569 = 43.95g

g = 8.12%
Answer #2

ANSWER :


Given :

r = 15.9% = 0.159 

P0 = 41 ($)

D0 = 2.95 ($)


Let dividend growth rate be g.. 

So, 

D1 = 2.95(1 + g) 

Now,

r = D1/P0 + g

=> 0.159 = 2.95(1 + g) / 41 + g 

=> 0.159 = 2.95/41 + g (2.95/41 + 1) 

=> 0.159 = 0.072 + 1.072 g

=> g = (0.159 - 0.072) / 1.072

=> g = 0.08116

=> g = 8.116 % = 8.12 % approx.


So,


8.12% is the dividend growth rate. (ANSWER)


answered by: Tulsiram Garg